# A systematic study of AGN feedback in a disk galaxy II: MACER prediction of X-ray surface brightness profile and comparison with eROSITA observations

### 23 Mar 2026 v1

@@ Abstract
Recently, we have performed a systematic study of AGN feedback in a disk galaxy within the MACER framework. Various model predictions, including the AGN duty cycle, the correlation between black hole accretion rates and star formation rates, and the (cold) gas fraction, have been compared with observations and will be presented in a series of papers. As the second paper in this series, without adjusting any model parameters, we directly use the simulation data introduced in Paper I to compute the predicted X-ray surface brightness profile and compare it with eROSITA observations of circumgalactic medium (CGM) emission around galaxies, which provide important constraints on AGN feedback models. For this comparison, we adopt two stacked eROSITA radial profiles of X-ray surface brightness: (1) distant galaxies with $\log(M_{*}/M_{\odot})=10.5-11.0$ at $z\approx 0.02-0.10$ from Zhang et al. (2024), and (2) nearby $L^{*}$ galaxies within 50 Mpc from He and Li (2026). We find that the average simulated profile over time is in good agreement with the stacked measurements of Zhang et al. (2024) over a broad radial range (out to $\sim 100\,\mathrm{kpc}$). Our model predictions also match the results of He and Li (2026) at projected radii from $\sim 20\,\mathrm{kpc}$ to $120\,\mathrm{kpc}$. Overall, the consistency between our simulations and the eROSITA data indicates that the X-ray emission detected by eROSITA is predominantly thermal in origin, rather than nonthermal, as supported by the spectral analysis presented by He and Li (2026).
@@ I Introduction
Several studies have compared the predictions from numerical simulations of AGN feedback and galaxy evolution with the observed X-ray surface brightness profiles (Shreeram et al., 2025; Lau et al., 2025; Grayson et al., 2025; Zhang et al., 2025c). Shreeram et al. (2025) used IllustrisTNG-based forward modeling to interpret eROSITA stacked X-ray profiles of Milky Way-mass galaxies, separating the contributions from hot CGM gas, satellites, and unresolved point sources. They found that the inferred CGM X-ray emission is strongly dependent on the underlying halo mass distribution, with models favoring $M_{200m}\approx 3.5\times 10^{12}M_{\odot}$ providing the best agreement with observations of Zhang et al. (2024) (hereafter Z24). Using the Cosmology and Astrophysics with MachinE Learning Simulations (CAMELS) suite, which consists of a large number of cosmological hydrodynamical simulations with varying cosmological parameters and baryonic feedback uniquely suited for studying feedback physics at the CGM, Lau et al. (2025) investigates how stellar and AGN feedback affect the X-ray properties of the hot CGM. They find that stronger feedback is needed than that currently implemented in the IllustrisTNG, SIMBA, and Astrid simulations to match the observed CGM properties. However, they emphasize that adopting these enhanced feedback parameters causes deviations in the stellar mass-halo mass relations from observational constraints below the group-mass scale.
In this work, we calculate the predicted X-ray emission of the CGM of a MW-mass galaxy based on the numerical simulation results presented in Zou et al. (2026), which are based on the MACER framework (Yuan et al., 2018), and compare the prediction with eROSITA observations. Since the simulations of Zou et al. (2026) are performed for a single galaxy, we approximate a galaxy ensemble with a time ensemble; that is, we stack outputs from different time steps of the same galaxy to mimic the effect of stacking different galaxies. Although real galaxies exhibit diversity in their initial conditions, dark matter halo properties, bulge-to-disk ratios, and other structural parameters, we argue that this approximation is reasonable for our purposes. For the X-ray surface brightness profile, the quantities of direct relevance are the gas density and temperature profiles, while the effects of other galaxy properties enter mainly through their influence on these thermodynamic quantities. Given that the diversity of these structural properties in the observed galaxy sample is limited, we expect galaxy mass and, in particular, AGN feedback to play a more important role in determining the X-ray surface brightness profile. For example, at different epochs, AGN feedback can change the gas content of a galaxy by as much as three orders of magnitude, which would translate into an approximately six-order-of-magnitude variation in the X-ray surface brightness, as we will show in the Results section below. In this sense, our treatment should capture the zeroth-order behavior of the observational stacks, although this caveat should be kept in mind when interpreting the comparison.
@@
The paper is organized as follows. In Section II, we briefly provide an overview of the main observational results, including the results from Z24, and He and Li (2026), in which a different $L_{*}$ sample is used. We describe the Zou et al. (2026) simulation in Section III. In Section IV.1, we calculate the surface brightness profile of the CGM predicted by our Fiducial model presented in Zou et al. (2026) and compare the results with observations. The effects of different initial conditions of our simulation are presented and discussed in Section IV.2. To quantitatively understand the effect of AGN feedback on controlling the CGM surface brightness profile, we have calculated the prediction of a model without AGN feedback and the results are presented in Section IV.3. We summarize our results in Section V.
@@
eROSITA is a highly sensitive soft X-ray telescope optimized for the $0.3\text{--}2.3\,\mathrm{keV}$ band, making it well suited for detecting emission from the hot circumgalactic medium (van Gelder et al., 2024; Predehl, P. et al., 2021). Z24 uses X-ray data from the first four SRG/eROSITA All-Sky Surveys to measure the extended X-ray emission from the hot CGM. Their analysis is based on a stacking measurement of 85,222 central galaxies in the Milky Way-mass bin ($\log(M_{*}/M_{\odot})=10.5\text{--}11.0$; $M_{*,\rm med}=5.5\times 10^{10}\,M_{\odot}$, i.e. $\log(M_{*,\rm med}/M_{\odot})\simeq 10.74$; Table 3 of Z24), drawn from an approximately volume-limited sample ($0.02<z<0.10$, $z_{\rm med}=0.08$) selected from the SDSS DR7 spectroscopic survey.
The central galaxy sample is constructed to achieve high completeness while reducing contamination from satellite galaxies. An important consideration in the analysis is the angular resolution of eROSITA, characterized by a point-spread function (PSF) of $\sim 30^{\prime\prime}$ (van Gelder et al., 2024). This PSF sets a practical limit on the redshift range over which the extended hot CGM can be spatially resolved and, therefore, informs both the galaxy sample selection and the stacking strategy.
@@ III.1 The MACER framework
The simulations are performed within the MACER framework, which is a model designed to simulate the evolution of an individual galaxy, with a particular focus on the role of AGN feedback. The physics of MACER is described in detail in Yuan et al. (2018). In the present paper, we use the two-dimensional version from Yuan et al. (2018). The model self-consistently follows the interplay among a supermassive black hole, a dark matter halo, a stellar bulge, a stellar disk, a rotationally supported gaseous disk, and cosmological inflow. Key physical processes, such as AGN and stellar feedback, star formation, and radiative cooling, are all incorporated.
Compared to large-scale cosmological simulations, MACER not only achieves significantly higher spatial resolution but also features a meticulous treatment of black hole accretion and feedback. The inner boundary of the simulation is set to be smaller than the Bondi radius of the accretion flow. By combining the mass flux at the inner boundary with the theory of black hole accretion, we can reliably calculate the black hole accretion rate at the event horizon, which determines the power of the AGN and is therefore essential for modeling AGN feedback. Once the mass accretion rate is obtained, the AGN outputs—namely radiation, wind, and jet—are prescribed based on accretion theory. Specifically, accretion is classified into cold and hot modes according to the accretion rate. The properties of wind and jet (e.g., velocity and mass flux) in the hot mode are derived directly from small-scale GRMHD simulations of black hole accretion and jet formation (e.g., Yang et al., 2021), while those in the cold mode are informed directly by observations (e.g., Gofford et al., 2015). These outputs are then injected at the inner boundary of the simulation domain, enabling a self-consistent calculation of their momentum and energy feedback on the ISM. This approach offers substantially greater precision than the phenomenological approach commonly adopted in cosmological simulations, where a fixed fraction of AGN power is often deposited into an assumed region around the AGN.
@@ III.2 Setup of the numerical simulation of Paper I
Figure 1: Simulated versus observed CGM X-ray surface-brightness profiles as a function of projected radius. Left: Purple points (with error bars) show the stacked profile of Milky Way-mass galaxies from Z24 (median distance $\sim$320 Mpc). Gold curves and shaded bands show predictions of the Fiducial model from Paper I for three gas metallicities: dashed, solid, and dotted lines correspond to $Z=0.02\,Z_{\odot}$, $0.3\,Z_{\odot}$, and $0.5\,Z_{\odot}$, respectively. For $Z=0.3\,Z_{\odot}$, the solid curve denotes the mean profile, and the shaded regions enclose the central 10%, 50%, and 90% percentile ranges. Right: Black squares show the stacked profile from He and Li (2026); median distance $\sim$36 Mpc. Because their count-rate–to–flux conversion assumes solar metallicity, we show the simulation only for $Z=1.0\,Z_{\odot}$. The green solid curve is the mean Fiducial prediction, with shaded regions indicating the central 10%, 50%, and 90% percentile ranges. The simulated profile is convolved with the same angular PSF as in the left panel and rescaled to the physical scale appropriate for the closer median distance.
Figure 2: Comparison between simulated and observed stacked X-ray surface brightness profiles of the CGM around Milky Way-mass galaxies. Left: Purple points with error bars show the stacked X-ray surface brightness profile from Z24, whose observational sample has a median distance of $\sim$320 Mpc. The simulated profiles are convolved with the same angular PSF, corresponding to this median distance. Right: Black squares show the stacked CGM X-ray surface brightness profile from He & Li (2025), with a median sample distance of $\sim$36 Mpc. The simulated profiles are convolved with the same angular PSF, but rescaled to the closer median distance. In both panels, the green solid line shows the mean profile of our Fiducial model, while the green dashed line shows the mean profile of the Fiducial_3 model, in which the initial CGM gas mass is three times that of the Fiducial model.
@@ IV.1 The Fiducial Model
The X-ray emissivity profile, $\epsilon_{\mathrm{X}}(r)$, is calculated using the simulation data as described below. For each radial bin $r$, the local X-ray volume emissivity is computed as $\epsilon_{\mathrm{X}}(r)=\Lambda(T,Z)n_{\mathrm{e}}n_{\mathrm{H}}$, where $n_{\mathrm{e}}$ and $n_{\mathrm{H}}$ are the electron and hydrogen number densities, $T$ is the gas temperature, $Z$ is the metallicity, and $\Lambda(T,Z)$ is the temperature- and metallicity-dependent X-ray cooling function (calculated using the APEC model). To enable direct comparison with eROSITA observations, the intrinsic 3D emissivity profile (which is obtained by revolving the 2D simulation data around the azimuthal ($\phi$) axis under the assumption of spherical symmetry) is first projected along the line of sight to generate a 2D surface brightness map $S_{\mathrm{X,intrinsic}}(x,y)$. This ideal map is then convolved with the instrumental PSF appropriate for the specific observatory and instrument configuration to produce the observable surface brightness $S_{\mathrm{X,obs}}(x,y)$.
We construct the mean radial profile in the simulations following the same pipeline adopted in the observations (Z24). Observationally, the mean X-ray surface brightness profile of a galaxy sample is obtained by summing the X-ray emission from individual galaxies without applying any additional weighting, and then measuring a radial profile from the stacked image. To mimic this stacking procedure in our simulations, we compute the X-ray surface brightness profile for each snapshot and then average the resulting profiles. For our simulated sample, these two averaging approaches are numerically equivalent. In doing so, we exclude simulation snapshots in which the AGN luminosity is significantly higher than that characteristic of the observational comparison samples.
@@
Figure 1 presents a comparison between the simulated and observed CGM X-ray surface brightness profiles. As our simulations follow the evolution of a single galaxy, we construct an average radial profile by time-averaging over the full set of simulation snapshots, which we interpret as representing an ensemble of galaxies. We exclude the interval from 7.8 to 8.2 Gyr in the Paper I simulations to avoid epochs dominated by a luminous quasar-mode AGN phase. This exclusion is well motivated, as the AGN in both observational samples considered here—Z24 and He and Li (2026)—generally exhibit modest luminosities rather than sustained high-accretion states. For a fair comparison with the observations, the simulated profiles are convolved with the same angular point-spread function (PSF), with angular scales converted to physical units using the median distance of each observational sample.
We focus our quantitative comparison on projected radii beyond $\sim 20\,\mathrm{kpc}$, where uncertainties associated with modeling unresolved eROSITA point sources in the galactic disk are minimized, as Z24 provides only weak constraints on the X-ray emission within $\sim 20\,\mathrm{kpc}$.
A key result is that the mean simulated profiles, averaged over metallicities $Z=0.02\,Z_{\odot}$, $0.3\,Z_{\odot}$, and $0.5\,Z_{\odot}$, are in good agreement with the stacked measurements of Z24 across a wide range of radii, out to $\sim 100\,\mathrm{kpc}$. In the case of He and Li (2026), the simulations likewise reproduce the observed surface brightness profile well at radii beyond $\sim 20\,\mathrm{kpc}$, consistent with our earlier findings. We further note that the distributions around the mean—quantified by the central 10%, 50% and 90% percentile ranges—span a relatively broad dynamic range, indicating substantial intrinsic temporal variability in the CGM X-ray emission. As a result, the time-averaged mean is weighted toward epochs of elevated surface brightness, which are typically associated with phases of enhanced AGN activity.
@@ IV.2 Predicting the CGM mass: the role of initial conditions
In observations, the total CGM mass still exhibits substantial uncertainty, mainly due to uncertainty in the mass of the hot phase, driven by the degeneracy between density, temperature, and metallicity (e.g., Bregman et al., 2018; Chen and Zahedy, 2026; Zhang et al., 2025b). In the Fiducial model of Paper I, the initial mass of the CGM is set to $2\times 10^{10}M_{\odot}$. In this model, the total mass of the CGM is regulated by several processes, including mass loss due to star formation and galactic outflows and mass increase due to stellar winds and cosmological inflow. The time-averaged total mass of the CGM within 400 kpc during the galaxy’s evolution, obtained from the Paper I simulations, is $\sim 5\times 10^{10}M_{\odot}$. The CGM mass exhibits considerable temporal variation, with the interquartile range (25th to 75th percentiles) spanning from $5.2\times 10^{9}$ to $9.3\times 10^{10}M_{\odot}$.
@@
The time-averaged total CGM gas mass predicted by the Fiducial_3 model is $3\times 10^{11}\,M_{\odot}$, approximately six times larger than that obtained in the Fiducial model of Paper I. This implies that the mean total CGM mass is not simply proportional to the initial gas mass, likely due to complex feedback interplay; for example, in the higher-initial-mass case, more gas can cool and condense into cold clumps that are less susceptible to feedback, allowing a larger fraction to remain in the CGM. Such a large total CGM mass leads to a severe overprediction of the X-ray surface brightness, thereby providing strong support for the robustness of our fiducial estimate of a typical total CGM mass of $\sim 5\times 10^{10}\,M_{\odot}$. This conclusion is consistent with recent stacked X-ray analyzes of Milky Way–mass galaxies, which indicate that CGM models with total gas masses approaching a few $\times 10^{11}\,M_{\odot}$ would generally exceed the observed X-ray emission unless extreme assumptions about the thermal structure or metallicity are adopted (Zhang et al., 2025a).
@@
Finally, in addition to our prediction for the representative CGM mass, the left panel of Figure 3 presents the temporal evolution of the total gas mass enclosed within $30\text{--}250\,\mathrm{kpc}$, facilitating a more direct comparison with existing observational constraints.
Figure 4: Comparison of simulated and observed CGM X-ray surface brightness profiles. Left: Purple points with error bars show the stacked X-ray surface brightness profile of Milky Way-mass galaxies from Z24. Solid gold and dashed blue curves represent the mean predictions of the Fiducial and noAGN models, respectively; shaded bands mark the central 10% percentile ranges. Right: Black squares show the stacked CGM X-ray surface brightness profile from He and Li (2026). Solid green and dashed blue curves represent the mean predictions of the Fiducial and noAGN models, respectively, with shading again indicating the central 10% percentile ranges. The comparison underscores the significant role of AGN feedback in shaping the CGM X-ray emission.
Figure 5: Time-averaged radial profiles of CGM gas properties. Left: Gas number density profile. Middle: Gas temperature profile. Right: Gas entropy profile. In all panels, the green solid curve corresponds to the Fiducial model, while the blue dashed curve represents the noAGN model.

### 01 May 2026 v2

# A systematic study of AGN feedback in a disk galaxy using MACER II: predictions of X-ray surface brightness profiles and comparison with eROSITA observations

@@ Abstract
Recently, we have performed a systematic study of AGN feedback in a disk galaxy within the MACER framework. Various model predictions, including the AGN duty cycle, the correlation between black hole accretion rates and star formation rates, and the (cold) gas fraction, have been compared with observations and will be presented in a series of papers. As the second paper in this series, without adjusting any model parameters, we directly use the simulation data introduced in Paper I to compute the predicted X-ray surface brightness profile and compare it with eROSITA observations of circumgalactic medium (CGM) emission around galaxies, which provide important constraints on AGN feedback models. For this comparison, we adopt two stacked eROSITA radial profiles of X-ray surface brightness: (1) distant galaxies with $\log(M_{*}/M_{\odot})=10.5-11.0$ at $z\approx 0.02-0.10$ from Zhang et al. (2024), and (2) nearby $L^{*}$ galaxies within 50 Mpc from He and Li (2026). We find that the average simulated profile over time is in good agreement with the stacked measurements of Zhang et al. (2024) over a broad radial range (out to $\sim 100\,\mathrm{kpc}$). Our model predictions also match the results of He and Li (2026) at projected radii from $\sim 20\,\mathrm{kpc}$ to $120\,\mathrm{kpc}$. While our simulations, which predict only thermal emission, are consistent with these recent X-ray observations, the limitations in our current model mean that this agreement does not preclude a potential contribution from non-thermal emission, e.g., from an extended halo of cosmic rays.
@@ I Introduction
Several studies have compared the predictions from numerical simulations of AGN feedback and galaxy evolution with the observed X-ray surface brightness profiles (Shreeram et al., 2025; Lau et al., 2025; Grayson et al., 2025; Zhang et al., 2025b). Shreeram et al. (2025) used IllustrisTNG-based forward modeling to interpret eROSITA stacked X-ray profiles of Milky Way-mass galaxies, separating the contributions from hot CGM gas, satellites, and unresolved point sources. They found that the inferred CGM X-ray emission is strongly dependent on the underlying halo mass distribution, with models favoring $M_{200m}\approx 3.5\times 10^{12}M_{\odot}$ providing the best agreement with observations of Zhang et al. (2024) (hereafter Z24). Using the Cosmology and Astrophysics with MachinE Learning Simulations (CAMELS) suite, which consists of a large number of cosmological hydrodynamical simulations with varying cosmological parameters and baryonic feedback uniquely suited for studying feedback physics at the CGM, Lau et al. (2025) investigates how stellar and AGN feedback affect the X-ray properties of the hot CGM. They find that stronger feedback is needed than that currently implemented in the IllustrisTNG, SIMBA, and Astrid simulations to match the observed CGM properties. However, they emphasize that adopting these enhanced feedback parameters causes deviations in the stellar mass-halo mass relations from observational constraints below the group-mass scale.
@@
In this work, we calculate the predicted X-ray emission of the CGM of a MW-mass galaxy based on the numerical simulation results presented in Zou et al. (2026), which are based on the MACER framework (Yuan et al., 2018), and compare the prediction with eROSITA observations. The paper is organized as follows. In Section II, we briefly provide an overview of the main observational results, including the results from Z24, and He and Li (2026), in which a different $L_{*}$ sample is used. We describe the Zou et al. (2026) simulation in Section III. In Section IV.1, we calculate the surface brightness profile of the CGM predicted by our Fiducial model presented in Zou et al. (2026) and compare the results with observations. The effects of different initial conditions of our simulation are presented and discussed in Section IV.2. To quantitatively understand the effect of AGN feedback on controlling the CGM surface brightness profile, we have calculated the prediction of a model without AGN feedback and the results are presented in Section IV.3. We summarize our results in Section V.
@@
eROSITA is a highly sensitive soft X-ray telescope optimized for the $0.3\text{--}2.3\,\mathrm{keV}$ band, making it well suited for detecting emission from the hot circumgalactic medium(Merloni, A. et al., 2024; Predehl, P. et al., 2021). Z24 uses X-ray data from the first four SRG/eROSITA All-Sky Surveys to measure the extended X-ray emission from the hot CGM. Their analysis is based on a stacking measurement of 85,222 central galaxies in the Milky Way-mass bin ($\log(M_{*}/M_{\odot})=10.5\text{--}11.0$; $M_{*,\rm med}=5.5\times 10^{10}\,M_{\odot}$, i.e. $\log(M_{*,\rm med}/M_{\odot})\simeq 10.74$; Table 3 of Z24), drawn from an approximately volume-limited sample ($0.02<z<0.10$, $z_{\rm med}=0.08$) selected from the SDSS DR7 spectroscopic survey.
The central galaxy sample is constructed to achieve high completeness while reducing contamination from satellite galaxies. An important consideration in the analysis is the angular resolution of eROSITA, characterized by a point-spread function (PSF) of $\sim 30^{\prime\prime}$ (Merloni, A. et al., 2024). This PSF sets a practical limit on the redshift range over which the extended hot CGM can be spatially resolved and, therefore, informs both the galaxy sample selection and the stacking strategy.
@@ III.1 The MACER framework
The simulations are performed within the MACER framework, which is a model designed to simulate the evolution of an individual galaxy, with a particular focus on the role of AGN feedback. The physics of MACER is described in detail in Yuan et al. (2018) [^1] Recently, the AGN-related physics introduced in MACER has been incorporated into the cosmological simulation IllustrisTNG (Zhu et al., 2026).. The model self-consistently follows the interplay among a supermassive black hole, a dark matter halo, a stellar bulge, a stellar disk, a rotationally supported gaseous disk, and cosmological inflow. Key physical processes, such as AGN and stellar feedback, star formation, and radiative cooling, are all incorporated.
Compared to large-scale cosmological simulations, the MACER model achieves significantly higher spatial resolution while providing a refined treatment of black hole accretion and feedback. While a comprehensive comparison is presented in He et al. (2025), we highlight several key distinctions below. The inner boundary of the simulation is set to be smaller than the Bondi radius of the accretion flow. By combining the mass flux at the inner boundary with the theory of black hole accretion, we can reliably calculate the black hole accretion rate at the event horizon, which determines the power of the AGN and is therefore essential for modeling AGN feedback. Once the mass accretion rate is obtained, the AGN outputs—namely radiation, wind, and jet—are prescribed based on accretion theory. Specifically, accretion is classified into cold and hot modes according to the accretion rate. The properties of wind and jet (e.g., velocity and mass flux) in the hot mode are derived directly from small-scale GRMHD simulations of black hole accretion and jet formation (e.g., Yang et al., 2021), while those in the cold mode are informed directly by observations (e.g., Gofford et al., 2015). These outputs are then injected at the inner boundary of the simulation domain, enabling a self-consistent calculation of their momentum and energy feedback on the ISM. This approach offers substantially greater precision than the phenomenological approach commonly adopted in cosmological simulations, where a fixed fraction of AGN power is often deposited into an assumed region around the AGN.
@@ III.2 Setup of the numerical simulation of Paper I
Figure 1: Simulated versus observed CGM X-ray surface-brightness profiles as a function of projected radius. Left: Purple points (with error bars) show the stacked profile of Milky Way-mass galaxies from Z24 (median distance $\sim$320 Mpc); the purple shaded region indicates the $1\sigma$ uncertainty of the best-fit $\beta$ model. Gold curves and shaded bands show predictions of the Fiducial model from Paper I for three gas metallicities: dashed, solid, and dotted lines correspond to $Z=0.02\,Z_{\odot}$, $0.3\,Z_{\odot}$, and $0.5\,Z_{\odot}$, respectively. For $Z=0.3\,Z_{\odot}$, the solid curve denotes the mean profile, and the shaded regions enclose the central 10%, 50%, and 90% percentile ranges. Right: Black squares show the stacked profile from He and Li (2026); median distance $\sim$36 Mpc. Because their count-rate–to–flux conversion assumes solar metallicity, we show the simulation only for $Z=1.0\,Z_{\odot}$. The green solid curve is the mean Fiducial prediction, with shaded regions indicating the central 10%, 50%, and 90% percentile ranges. The simulated profile is convolved with the same angular PSF as in the left panel and rescaled to the physical scale appropriate for the closer median distance.
Figure 2: Comparison between simulated and observed stacked X-ray surface brightness profiles of the CGM around Milky Way-mass galaxies. Left: Purple points with error bars show the stacked X-ray surface brightness profile from Z24, whose observational sample has a median distance of $\sim$320 Mpc; the purple shaded region indicates the $1\sigma$ uncertainty of the best-fit $\beta$ model. The simulated profiles are convolved with the same angular PSF, corresponding to this median distance. Right: Black squares show the stacked CGM X-ray surface brightness profile from He & Li (2025), with a median sample distance of $\sim$36 Mpc. The simulated profiles are convolved with the same angular PSF, but rescaled to the closer median distance. In both panels, the green solid line shows the mean profile of our Fiducial model, while the blue line shows the mean profile of the Fiducial_3 model, in which the initial CGM gas mass is three times that of the Fiducial model.
@@ IV.1 The Fiducial Model
The X-ray emissivity profile, $\epsilon_{\mathrm{X}}(r)$, is calculated using the simulation data in the 0.5–2 $\mathrm{keV}$ band as described below. For each radial bin $r$, the local X-ray volume emissivity is computed as $\epsilon_{\mathrm{X}}(r)=\Lambda(T,Z)\,n_{\mathrm{e}}n_{\mathrm{H}}$, where $n_{\mathrm{e}}$ and $n_{\mathrm{H}}$ are the electron and hydrogen number densities, $T$ is the gas temperature, $Z$ is the metallicity, and $\Lambda(T,Z)$ is the temperature- and metallicity-dependent X-ray cooling function (calculated using the APEC model). To enable direct comparison with eROSITA observations, the intrinsic 3D emissivity profile (which is obtained by revolving the 2D simulation data around the azimuthal ($\phi$) axis under the assumption of axial symmetry) is first projected along the line of sight to generate a 2D surface brightness map $S_{\mathrm{X,intrinsic}}(x,y)$. This ideal map is then convolved with the instrumental PSF appropriate for the specific observatory and instrument configuration to produce the observable surface brightness $S_{\mathrm{X,obs}}(x,y)$.
We construct the mean radial profile in the simulations following the same pipeline adopted in the observations (Z24). Observationally, the mean X-ray surface brightness profile of a galaxy sample is obtained by summing the X-ray emission from individual galaxies without applying any additional weighting and then measuring a radial profile from the stacked image. To mimic this stacking procedure in our simulations, we compute the X-ray surface brightness profile for each snapshot and then average the resulting profiles. For our simulated sample, these two averaging approaches are numerically equivalent. In doing so, we exclude simulation snapshots in which the AGN luminosity is significantly higher than that characteristic of the observational comparison samples.
@@
Figure 1 presents a comparison between the simulated and observed CGM X-ray surface brightness profiles. We exclude the interval from 7.8 to 8.2 Gyr in the Paper I simulations to avoid epochs dominated by a luminous quasar-mode AGN phase. This exclusion is well motivated, as the AGN in both observational samples considered here—Z24 and He and Li (2026)—generally exhibit modest luminosities rather than sustained high-accretion states. For a fair comparison with the observations, the simulated profiles are convolved with the same angular point-spread function (PSF), with angular scales converted to physical units using the median distance of each observational sample.
We focus our quantitative comparison on projected radii beyond $\sim 20\,\mathrm{kpc}$, where uncertainties associated with modeling unresolved eROSITA point sources in the galactic disk are minimized, as Z24 provides only weak constraints on the X-ray emission within $\sim 20\,\mathrm{kpc}$. A key result is that the mean simulated profiles, averaged over metallicities $Z=0.02\,Z_{\odot}$, $0.3\,Z_{\odot}$, and $0.5\,Z_{\odot}$, are in good agreement with the stacked measurements of Z24 across a wide range of radii, out to $\sim 100\,\mathrm{kpc}$. In the case of He and Li (2026), the simulations likewise reproduce the observed surface brightness profile well at radii beyond $\sim 20\,\mathrm{kpc}$. We further note that the distributions around the mean—quantified by the central 10%, 50% and 90% percentile ranges—span a relatively broad dynamic range, indicating substantial intrinsic temporal variability in the CGM X-ray emission. As a result, the time-averaged mean is weighted toward epochs of elevated surface brightness, which are typically associated with phases of enhanced AGN activity.
@@ IV.2 Predicting the CGM mass: the role of initial conditions
In observations, the total CGM mass still exhibits substantial uncertainty, mainly due to uncertainty in the mass of the hot phase, driven by the degeneracy between density, temperature, and metallicity (e.g., Bregman et al., 2018; Chen and Zahedy, 2026; Zhang et al., 2025a). In the Fiducial model of Paper I, the initial mass of the CGM is set to $2\times 10^{10}M_{\odot}$. In this model, the total mass of the CGM is regulated by several processes, including mass loss due to star formation and galactic outflows and mass increase due to stellar winds and cosmological inflow. The time-averaged total mass of the CGM within 400 kpc during the galaxy’s evolution, obtained from the Paper I simulations, is $\sim 5\times 10^{10}M_{\odot}$. The CGM mass exhibits considerable temporal variation, with the interquartile range (25th to 75th percentiles) spanning from $5.2\times 10^{9}$ to $9.3\times 10^{10}M_{\odot}$.
@@
The time-averaged total CGM gas mass predicted by the Fiducial_3 model is $3\times 10^{11}\,M_{\odot}$, approximately six times larger than that obtained in the Fiducial model of Paper I. This implies that the mean total CGM mass is not simply proportional to the initial gas mass, likely due to complex feedback interplay; for example, in the higher-initial-mass case, more gas can cool and condense into cold clumps that are less susceptible to feedback, allowing a larger fraction to remain in the CGM. Such a large total CGM mass leads to a severe overprediction of the X-ray surface brightness, thereby providing strong support for the robustness of our fiducial estimate of a typical total CGM mass of $\sim 5\times 10^{10}\,M_{\odot}$. This conclusion is consistent with recent stacked X-ray analyzes of Milky Way–mass galaxies, which indicate that CGM models with total gas masses approaching a few $\times 10^{11}\,M_{\odot}$ would generally exceed the observed X-ray emission unless extreme assumptions about the thermal structure or metallicity are adopted (Zhang, Yi et al., 2026).
@@
Finally, in addition to our prediction for the representative CGM mass, the left panel of Figure 3 presents the temporal evolution of the total gas mass enclosed within $30\text{--}250\,\mathrm{kpc}$. This range is chosen to facilitate a more direct comparison with existing observational constraints: while the hot CGM can extend to larger radii, observational measurements of CGM gas mass in MW-mass galaxies are typically reported within the virial radius, which corresponds to $\sim$200–300 kpc for this mass regime. Restricting the integration to 30–250 kpc therefore places our model predictions on a more comparable footing with such constraints.
Figure 4: Comparison of simulated and observed CGM X-ray surface brightness profiles. Left: Purple points with error bars show the stacked X-ray surface brightness profile of Milky Way-mass galaxies from Z24; the purple shaded region indicates the $1\sigma$ uncertainty of the best-fit $\beta$ model. Solid gold and dashed blue curves represent the mean predictions of the Fiducial and noAGN models, respectively; shaded bands mark the central 10% percentile ranges. Right: Black squares show the stacked CGM X-ray surface brightness profile from He and Li (2026). Solid green and dashed blue curves represent the mean predictions of the Fiducial and noAGN models, respectively, with shading again indicating the central 10% percentile ranges. The comparison underscores the significant role of AGN feedback in shaping the CGM X-ray emission.
Figure 5: Time-averaged radial profiles of CGM gas properties. The gas number density profile is volume weighted, while the gas temperature and gas entropy profiles are gas mass weighted. Left: Gas number density profile. Middle: Gas temperature profile. Right: Gas entropy profile. In all panels, the green solid curve corresponds to the Fiducial model, while the blue dashed curve represents the noAGN model.